# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 1 Introduction
In the evolving field of medical image analysis, the dynamic nature of healthcare data poses a critical challenge for the generalizability of the machine learning/deep learning models to new data/domains Miotto et al. (2017); Zhou et al. (2021a). The data-driven approaches have challenges due to the limited availability and accessibility of sufficiently large and diverse medical data for training Miotto et al. (2017); Guan and Liu (2022); Zhou et al. (2021a); Chauhan and Goyal (2020, 2021). Additionally, the source variability, scanner manufacturer, imaging protocol, and diverse population groups make the medical data heterogeneous, this adds a bias and discrepancies between training and testing datasets, thus leading to performance degradation Oliveira et al. (2023); Zhou et al. (2021a). In order to handle the model generalizability issues, domain adaptation methods become popular and aim to transfer knowledge from one domain to other unseen data sources or domains Lai et al. (2023a, b); Guan and Liu (2022); CW (2016); Becker et al. (2014); FEN (2023). However, domain adaptation poses unique challenges due to the sensitive and complex nature of healthcare data Azad et al. (2022). The most common associated issues are the limited availability of labeled medical data for training, heterogeneity in data sources contributing to significant domain shifts, clinical disparities and population variances, and inter-rater variabilities. Also, biases present in the source domain data can propagate to the target domain, so ensuring fair and unbiased predictions across diverse patient populations becomes a critical concern. Moreover, medical data encompasses various modalities, including imaging, electronic health records, and genomic data, making the task of adapting models to handle multimodal data and ensuring interoperability exceptionally complex.
@@
Continual learning (CL) - continuous adaptation to new information, has emerged as an important dimension in enhancing the performance and reliability of medical image analysis systems. Transfer learning is widely adopted in the medical domain to address challenges related to limited data availability Yu et al. (2022); Ghafoorian et al. (2017). It transfers knowledge gained from the source task to the target task to improve its learning or performance. The focus is on leveraging prior knowledge rather than retaining it. Unlike transfer learning, CL focuses on both the source domain and the target domain. A CL aims to retain knowledge from previously seen tasks while adapting to new tasks and avoiding catastrophic forgetting issues. CL models are employed for predictive analytics, especially in situations where clinical outcomes can be automatically obtained and incorporated into the algorithm. This capability enhances the model’s predictive power by learning from real-time patient data. CL methods can also be employed to utilize the multi-modality dataset for better interpretability and analysis. In recent years, an increasing number of CL methods have been explored and proposed in various subareas of computer vision tasks. Figure 1 presents the paper distributions based on different aspects of CL and shows the growth in exploration in this domain through the rising number of publications over the years. In this paper, we discuss the various aspects of CL, particularly considering its application and implications in the medical domain. Here, we aim to contribute to the ongoing discourse on adapting and improving machine learning models for sustained effectiveness in the dynamic healthcare landscape. By emphasizing CL, we recommend models that not only demonstrate robust performance initially but also possess the ability to evolve and improve over time. CL in medical image analysis represents a powerful approach to developing intelligent systems that can evolve, learn, and adapt to the complexities of healthcare, ultimately contributing to improved patient outcomes and enhanced clinical decision-making. We discuss the limitations and challenges of the existing methods and explore the methods/techniques that can be utilized for developing a robust algorithm. A thorough search of the existing body of literature reveals the groundbreaking nature of our work as the first comprehensive survey of CL techniques applied to medical image analysis. This scholarly endeavor aims not only to contribute novel insights but also to establish a foundational reference for researchers, offering a roadmap that can guide future exploration and incite scholarly interest in the academic community. The primary contributions of our academic pursuit are elucidated below:
@@
Over the past few decades, CL approaches in computer vision tasks have seen substantial progress, leading to numerous survey papers exploring deep CL models for computer vision tasks Qu et al. (2021); Wang et al. (2023a); De Lange and Tuytelaars (2021); Mai et al. (2022). While some of these surveys focus on specific applications, such as classifications De Lange and Tuytelaars (2021), others take a more general approach to concepts and practical perspectives Wang et al. (2023a). Notably, none of these surveys specifically addresses the applications of CL techniques in medical image analysis, leaving a significant gap in the literature. We believe that insights from successful CL models in vision can be beneficial for the medical community, guiding the retrospective analysis of past and future research directions in CL Verma et al. (2023); Lee and Lee (2020). CL has proven its potential in developing unified and sustainable deep models capable of handling new classes, tasks, and the evolving nature of data in non-stationary environments across various application areas. Our survey aims to bridge the gap by providing valuable insights that can assist medical researchers, including radiologists, in adopting up-to-date methodologies in their fields. In our survey, we analyze various sources of drifts in medical data, defining CL scenarios in medical images. We present a multi-perspective view of CL by categorizing techniques into rehearsal, regularization, architectural, and hybrid strategies-based methods. The discussion extends beyond applications, encompassing underlying working principles, challenges, and the imaging modality of the proposed methods. We emphasize how this additional information can aid researchers in consolidating literature across the spectrum. A concise overview of our paper is illustrated in Figure 6.
@@
To conduct a thorough literature search, we followed the same strategy presented in Azad et al.; Azad et al. (2023a; 2023b) and utilized DBLP, Google Scholar, and Arxiv Sanity Preserver, employing custom search queries to retrieve scholarly publications related to our topic-CL. Our search query was (continual learning | medical | sequence of tasks) (segmentation | classification | medical | lifelong learning). These platforms allowed us to filter results into categories such as peer-reviewed journal papers, conference or workshop proceedings, non-peer-reviewed papers, and preprints. We carefully filtered our search results through a two-step process: initial screening based on the title and abstract, followed by a subsequent screening of the full text. This approach was instrumental in eliminating false positives, ensuring that only relevant content was included in our study.
@@
All the abbreviations used in this manuscript along with their expansion are tabulated in Table 1. A background about various sources of drift in medical data, the catastrophic forgetting issue, CL, a pipeline for CL framework, and CL applications is provided in Section 2. Then a thorough description of various kinds of continual learning scenarios found in the medical domain is elaborated in Section 3. Further, we cover various categories of continual learning techniques and their applicability in the medical domain via Section 4. Section 5 provides details about the level of supervision required in different kinds of proposed frameworks. Section 6 provides comprehensive practical information such as the experimental setups, training process, and evaluation metrics for measuring the plasticity and stability of continual learning frameworks. Section 7 discusses the current challenges in the continual learning literature and future directions. Eventually, the last section provides the conclusion of the survey.
Table 1: Abbreviations and their expansion. Green rows signify abbreviations related to CL domains (scenarios and techniques), cyan corresponds to imaging modalities and techniques, while orange is associated with evaluation metrics.
@@
Acronym Full name
@@ 2.1 Medical data drifts
In clinical practices, the data distribution evolves over time, reflecting the dynamic nature of the healthcare environment Sahiner et al. (2023a); Lacson et al. (2022); Moreno-Torres et al. (2012). Inconsistencies in data collection procedures across different healthcare settings or institutions contribute to data drift. Also, the introduction of new medical imaging devices, diagnostic tools, and data acquisition techniques leads to a shift in the technological landscape Derakhshani et al. (2022); Lacson et al. (2022). Moreover, the advancement in medical research and the discovery of new diseases/treatment methods raise the understanding of healthcare. This new knowledge can influence the characteristics of medical data that cause shifts in the underlying distribution Lacson et al. (2022); Moreno-Torres et al. (2012). Also, data sources can have their dynamics and therefore are inherently non-constant. For instance, cardiac CT images are captured under time-varying factors such as breathing and heart rates. Non-homogeneous data is another challenge, as individual health differences among patients can vary over time due to factors like genetics, age, occupation, and lifestyle Lacson et al. (2022); Sahiner et al. (2023b). Additionally, variations in sample preparation or pre-processing methods contribute to distinctions among imaging datasets. In digital histopathology, differences in staining policies across labs introduce undesired stain variances in whole-slide images Nakagawa et al. (2023); Madabhushi and Lee (2016). The imaging solution may also influence the final digital visualization throughout the entire learning process over time. Nonlinear augmentation of computed radiography occurs at various degrees due to the differing physical and chemical properties of contrast mediums from different brands. Variables like sensor signal-to-noise ratio (SNR), customized parameter settings in imaging software, and storage-friendly distortions can all impact the quality of the resultant image. For instance, in digital histopathology, billion-pixel whole slide images (WSI) at a fixed magnification have seen variations in storage size, ranging from megabytes to gigabytes per image over the years, consequently enhancing the dataset quality in terms of micron-per-pixel (MPP) for continual learning tasks Nakagawa et al. (2023).
@@
The medical data drifts can have significant implications for the performance and reliability of machine learning and deep learning models Sahiner et al. (2023b). Traditional machine learning often relies heavily on static data and feature engineering, where human experts manually select the relevant features. In the case of data drift, these handcrafted features may become less informative, and the model may struggle to adapt to new patterns BAY (2022). More specifically, for both machine learning and deep learning, this issue is particularly prevalent in dynamic and non-stationary environments, where the statistical properties of the data evolve. Understanding the sources of drift, such as inconsistencies in data collection and technological advancements, is crucial for developing robust models Sahiner et al. (2023b). Proactive strategies, including regular model updates and continuous monitoring, are essential to ensure that machine learning models remain effective and reliable in navigating the evolving healthcare landscape. By addressing medical data drift and developing machine learning models that can adapt to the ever-changing nature of clinical data, ultimately enhancing patient care and outcomes.
@@ 2.2 Catastrophic forgetting
Throughout a lifetime, a human brain continuously acquires knowledge, and learning new concepts or tasks has no detrimental effect on previously learned ones. Instead, learning several closely related concepts even boosts the learning of all associated ones. In contrast, artificial neural networks, although inspired by the human brain, often suffer from ‘catastrophic forgetting’, a tendency to overwrite or forget the knowledge acquired in the past upon learning new concepts McCloskey and Cohen (1989); Ratcliff (1990). This can be attributed to the fact that the model entirely optimizes for the given dataset. In other words, a model with optimized weights for a task T1, when trained on a new task T2, will freely optimize the existing weights to meet the objectives in task T2, which may now no longer be optimal for the previous task T1. This can be a significant challenge, especially in scenarios where an AI system is expected to learn and adapt to a stream of tasks or datasets over time. Catastrophic forgetting in neural networks is an interesting phenomenon that has attracted lots of attention in recent research Goodfellow et al. (2013); Kumari et al. (2024).
@@ 2.3 Continual learning overview
Alternately, CL, also termed as ‘continuously learning’, ‘incremental learning’, ‘sequential learning’ or ‘lifelong learning’, has emerged as a promising solution in various fields to deal with the catastrophic forgetting issue De Lange and Tuytelaars (2021); Mai et al. (2022). It helps in efficiently leveraging existing knowledge and incorporating new information without the need for extensive retraining. The primary goal of CL is to develop techniques and strategies that allow a neural network to learn new tasks while retaining knowledge of previous tasks. In other words, it aims to enable the network to continually adapt to new information without completely erasing or degrading its performance on earlier tasks. Overall, CL helps to address the issue of catastrophic forgetting and minimizes the need for additional resources to store historical data. CL offers a range of strategies and methods, such as regularization (constraining weight update to avoid forgetting learnt concepts), rehearsal (partially using some form of old data to replay with current data), and architectural modifications (reserving or partitioning network for different tasks), to help neural networks remember and consolidate knowledge from past tasks. These strategies help prevent or reduce catastrophic forgetting and improve the generalization ability of the model.

### v4

@@ 1 Introduction
In the evolving field of medical image analysis, the dynamic nature of healthcare data poses a critical challenge for the generalizability of the machine learning/deep learning models to new data/domains Miotto et al. (2017); Zhou et al. (2021a). The data-driven approaches have challenges due to the limited availability and accessibility of sufficiently large and diverse medical data for training Miotto et al. (2017); Guan and Liu (2022); Zhou et al. (2021a); Chauhan and Goyal (2020, 2021). Additionally, the source variability due to different scanner manufacturers, staining and imaging protocols, slice thickness, different patient cohorts, etc., makes the medical data heterogeneous. This introduces bias and discrepancies between the training and test datasets if they originate from different data sources, thus leading to performance degradation Oliveira et al. (2023); Zhou et al. (2021a). In order to handle the model generalizability issues, domain adaptation methods become popular and aim to transfer knowledge from one domain to other unseen data sources or domains Lai et al. (2023a, b); Guan and Liu (2022); CW (2016); Becker et al. (2014); FEN (2023). However, domain adaptation poses unique challenges due to the sensitive and complex nature of healthcare data Azad et al. (2022). The most common associated issues are the limited availability of labeled medical data for training, heterogeneity in data sources contributing to significant domain shifts, clinical disparities and population variances, and inter-rater variabilities. Also, biases present in the source domain data can propagate to the target domain, so ensuring fair and unbiased predictions across diverse patient populations becomes a critical concern. Moreover, medical data encompasses various modalities, including imaging, electronic health records, and genomic data, making the task of adapting models to handle multimodal data and ensuring interoperability exceptionally complex. Further, the accessibility of source data may be limited to a short period of time or may be prohibited altogether due to strict privacy regulations in the medical domain  Thandiackal et al. (2024). Thus, domain adaptation approaches that require simultaneous availability of source and target data may not be feasible.
Another related learning paradigm, transfer learning, has been widely adopted in the medical domain to address challenges related to limited data availability Yu et al. (2022); Ghafoorian et al. (2017). It transfers knowledge gained from the source task to the target task to improve its learning or performance. Unlike domain adaptation, where only the data distribution changes, transfer learning covers changes in the feature space, label space, as well as in the data distribution of the source and target domain  Kouw and Loog (2018); Zhou et al. (2022); Guan and Liu (2022). In a small-scale medical disease classification dataset, it can be beneficial to include knowledge gained from a model trained on a large-scale labeled natural image dataset (ImageNet). The model performance on the medical disease dataset may be better as compared to training the same model solely on the medical disease dataset from scratch. However, at the same time the performance on the ImageNet dataset cannot be guaranteed by this model (which is also not intended in transfer learning). In transfer learning, the focus is on leveraging prior knowledge rather than retaining it, hence performance on the source data may be compromised. Generally, learning a new dataset with shifted distribution results in a sharp decrease in performance on the source dataset, also known as “catastrophic forgetting” of deep neural networks  McCloskey and Cohen (1989); Goodfellow et al. (2013). A detailed description of catastrophic forgetting is provided in Section 2.2.
@@
In the real world, sequential adaptation to more than one target domain, without necessitating the availability of source data can be desired. In this direction, Continual learning (CL) - continuous adaptation to new information, has emerged as an important dimension in enhancing the performance and reliability of medical image analysis systems. Unlike transfer learning, CL focuses on both the source domain and the target domain. A CL approach aims to retain knowledge from previously seen tasks while adapting to new tasks and avoiding catastrophic forgetting issues. CL models are employed for predictive analytics, especially in situations where clinical outcomes can be automatically obtained and incorporated into the algorithm Lee and Lee (2020). This capability enhances the model’s predictive power by learning from real-time patient data. CL methods can also be employed to utilize the multi-modality dataset for better interpretability and analysis. In recent years, an increasing number of CL methods have been explored and proposed in various subareas of computer vision tasks. Figure 3 presents the paper distributions based on different aspects of CL and shows the growth in exploration in medical domain through the rising number of publications over the years. In this paper, we discuss the various aspects of CL, particularly considering its application and implications in the medical domain. Here, we aim to contribute to the ongoing discourse on adapting and improving machine learning models for sustained effectiveness in the dynamic healthcare landscape. By emphasizing CL, we recommend models that not only demonstrate robust performance initially but also possess the ability to evolve and improve over time. CL in medical image analysis represents a powerful approach to developing intelligent systems that can evolve, learn, and adapt to the complexities of healthcare, ultimately contributing to improved patient outcomes and enhanced clinical decision-making. We discuss the limitations and challenges of the existing methods and explore the methods/techniques that can be utilized for developing a robust algorithm. A thorough search of the existing body of literature highlights the uniqueness of our work as the first comprehensive survey of CL techniques applied to medical image analysis. This scholarly endeavor aims not only to contribute novel insights but also to establish a foundational reference for researchers, offering a roadmap that can guide future exploration and incite scholarly interest in the academic community. The primary contributions of our academic pursuit are elucidated below:
@@
Over the past few decades, CL approaches in computer vision tasks have seen substantial progress, leading to numerous survey papers exploring deep CL models for computer vision tasks Qu et al. (2021); Wang et al. (2023a); De Lange et al. (2021); Mai et al. (2022). While some of these surveys focus on specific applications, such as classifications De Lange et al. (2021); Mai et al. (2022) others take a more general approach to evaluation policies Mundt et al. (2021) or concepts and practical perspectives Wang et al. (2023a). Notably, none of these surveys specifically addresses the applications of CL techniques in medical image analysis, leaving a significant gap in the literature. We believe that insights from successful CL models in vision can be beneficial for the medical community, guiding the retrospective analysis of past and future research directions in CL Verma et al. (2023); Lee and Lee (2020). CL has proven its potential in developing unified and sustainable deep models capable of handling new classes, tasks, and the evolving nature of data in non-stationary environments across various application areas. Our survey aims to bridge the gap by providing valuable insights that can assist medical researchers, including radiologists, in adopting up-to-date methodologies in their fields. In our survey, we analyze various sources of drifts in medical data, defining CL scenarios in medical images. We present a multi-perspective view of CL by categorizing techniques into rehearsal, regularization, architectural, and hybrid strategies-based methods. The discussion extends beyond applications, encompassing underlying working principles, challenges, and the imaging modality of the proposed methods. We emphasize how this additional information can aid researchers in consolidating literature across the spectrum. A concise overview of our paper is illustrated in Figure 6.
@@
To conduct a thorough literature search, we followed the same strategy presented in Azad et al.; Azad et al. (2023a; 2023b) and utilized DBLP, Google Scholar, and Arxiv Sanity Preserver, employing custom search queries to retrieve scholarly publications related to our topic-CL. Our search query was (continual learning | medical | sequence of tasks) (segmentation | classification | medical | lifelong learning). These platforms allowed us to filter results into categories such as peer-reviewed journal papers, conference or workshop proceedings, non-peer-reviewed papers, and preprints. We filtered our search results through a two-step process: first by screening titles and abstracts, and then by reviewing the full text based on specific criteria as follows: 1. Relevance to Continual Learning: Focus on significant contributions to continual learning or related areas such as lifelong learning, incremental learning, and online learning. 2. Publication Venue: Preference for papers published in reputable journals and conferences known for high-quality research. 3. Novelty and Contribution: Inclusion of papers presenting novel research, methodologies, or applications. 4. Experimental Rigor: Papers must include comprehensive experiments and results validating their claims. 5. Theoretical Foundation: Preference for papers with a strong theoretical foundation, including formal definitions and analysis.
@@
All the abbreviations used in this manuscript along with their expansion are tabulated in Table 1. A background about various sources of drift in medical data, the catastrophic forgetting issue, CL, a pipeline for CL framework, and CL applications is provided in Section 2. Then a thorough description of various kinds of continual learning scenarios explored in the medical domain is elaborated in Section 3. Further, we cover various categories of continual learning techniques and their applicability in the medical domain via Section 4. Section 5 provides details about the level of supervision required in different kinds of proposed frameworks. Section 6 provides comprehensive practical information such as the experimental setups, training process, and evaluation metrics for measuring the plasticity and stability of continual learning frameworks. Section 7 discusses the current challenges in the continual learning literature and future directions. Eventually, the last section provides the conclusion of the survey.
Table 1: Abbreviations and their expansion. Green rows signify abbreviations related to CL (scenarios and techniques), cyan corresponds to imaging modalities and techniques, while orange is associated with evaluation metrics.
@@
sEMG Surface electromyography
WSI Whole Slide Images
wCC weighted Correlation
PCC Pearson Correlation Coefficient
rMSE root Mean Square Error
FLOPS Floating Point Operations Per Second
SRC Spearman’s rank correlation coefficient
KRC Kendall rank correlation coefficient
AUPRC Area Under the Precision-Recall Curve
@@
TPR True Positive Rate
SNR Signal-to-Noise Ratio
@@ 2.1 Medical data drifts
In clinical practices, the data distribution evolves over time, reflecting the dynamic nature of the healthcare environment Sahiner et al. (2023a); Lacson et al. (2022); Moreno-Torres et al. (2012). Inconsistencies in data collection procedures across different healthcare settings or institutions contribute to data drift. Also, the introduction of new medical imaging devices, diagnostic tools, and data acquisition techniques leads to a shift in the technological landscape Derakhshani et al. (2022); Lacson et al. (2022). Moreover, the advancement in medical research and the discovery of new diseases/treatment methods raise the understanding of healthcare. This new knowledge can influence the characteristics of medical data that cause shifts in the underlying distribution Lacson et al. (2022); Moreno-Torres et al. (2012). Also, data sources can have their dynamics and, therefore, are inherently non-constant. For instance, cardiac CT images are captured under time-varying factors such as breathing and heart rates. Non-homogeneous data is another challenge, as individual health differences among patients can vary over time due to factors like genetics, age, occupation, and lifestyle Lacson et al. (2022); Sahiner et al. (2023b). Additionally, variations in sample preparation or pre-processing methods contribute to distinctions among imaging datasets. In digital histopathology, differences in staining policies across labs introduce undesired stain variances in whole-slide images Nakagawa et al. (2023); Madabhushi and Lee (2016). The imaging solution may also influence the final digital visualization throughout the entire learning process over time. Nonlinear augmentation of computed radiography occurs at various degrees due to the differing physical and chemical properties of contrast mediums from different brands. Variables like sensor signal-to-noise ratio (SNR), customized parameter settings in imaging software, and storage-friendly distortions can all impact the quality of the resultant image. For instance, in digital histopathology, billion-pixel whole slide images (WSI) at a fixed magnification have seen variations in storage size, ranging from megabytes to gigabytes per image over the years, consequently enhancing the dataset quality in terms of micron-per-pixel (MPP) for continual learning tasks Nakagawa et al. (2023).
@@
Data drifts can be broadly categorized as the covariate, label, and concept shift. We provide their explanation along with examples in Table 2. The medical data drifts can have significant implications for the performance and reliability of machine learning and deep learning models Sahiner et al. (2023b). Traditional machine learning often relies heavily on static data and feature engineering, where human experts manually select the relevant features. In the case of data drift, these handcrafted features may become less informative, and the model may struggle to adapt to new patterns BAY (2022). More specifically, for both machine learning and deep learning, this issue is particularly prevalent in dynamic and non-stationary environments, where the statistical properties of the data evolve. Understanding the sources of drift, such as inconsistencies in data collection and technological advancements, is crucial for developing robust models Sahiner et al. (2023b). Proactive strategies, including regular model updates and continuous monitoring, are essential to ensure that machine learning models remain effective and reliable in navigating the evolving healthcare landscape by addressing medical data drift and developing machine learning models that can adapt to the ever-changing nature of clinical data, ultimately enhancing patient care and outcomes.
Table 2: Data drift categorization
Drift type Detail with example
Covariate drift It refers to situations when input distribution $p(x)$ differs and conditional distribution $p(y|x)$ remains same between source and target data. Example: Change in WSI staining type induces covariate drift. The model trained on H&E stained breast cancer data may not perform well on the same tissue with CD8 staining.
Label drift It refers to situations when output distribution $p(y)$ changes, but the conditional distribution $p(x|y)$ remains same. Example: Inter-annotators difference may induce this kind of drift as some experts may be biased toward annotating a particular disease class. Thus, some classes may be undersampled or oversampled in target compared to the source data.
Concept drift It refers to situations when input and output distribution remains same but the conditional relation $p(y|X)$ differs between source and target data. Example: association of chest X-ray to COVID +ve class changed over time with new findings.
@@ 2.2 Catastrophic forgetting
Throughout a lifetime, a human brain continuously acquires knowledge, and learning new concepts or tasks has no detrimental effect on previously learned ones. Instead, learning several closely related concepts even boosts the learning of all associated ones. In contrast, artificial neural networks, although inspired by the human brain, often suffer from ’catastrophic forgetting’, a tendency to overwrite or forget the knowledge acquired in the past upon learning new concepts McCloskey and Cohen (1989); Ratcliff (1990). This can be attributed to the fact that the model entirely optimizes for the given dataset. In other words, a model with optimized weights for a task T1, when trained on a new task T2, will freely optimize the existing weights to meet the objectives in task T2, which may now no longer be optimal for the previous task T1. This can be a significant challenge, especially in scenarios where an AI system is expected to learn and adapt to a stream of tasks or datasets over time. Catastrophic forgetting in neural networks is an interesting phenomenon that has attracted lots of attention in recent research Goodfellow et al. (2013); Kumari et al. (2024). Medical data often come from different sources with varying imaging protocols, equipment, and patient populations. For example, MRI scans from different hospitals may have distinct characteristics, leading to domain shifts that can cause a model to forget previously learned features when introduced to new data. Also, medical datasets are often limited in size and can be highly imbalanced, with some conditions being much more common than others. This imbalance can exacerbate catastrophic forgetting, as the model may overly specialize in newly introduced, more frequent classes at the expense of older, less frequent ones. Addressing catastrophic forgetting is crucial for the development of reliable and effective medical AI systems. By implementing strategies such as regularization, rehearsal, generative replay, dynamic architectures, and domain adaptation, researchers can enhance the robustness of models against forgetting, ensuring consistent and accurate performance in medical applications. This is essential for maintaining diagnostic consistency, adapting to new medical knowledge, and providing trustworthy clinical decision support.
@@ 2.3 Continual learning overview
Alternately, CL, also termed as ‘continuously learning’, ‘incremental learning’, ‘sequential learning’ or ‘lifelong learning’, has emerged as a promising solution in various fields to deal with the catastrophic forgetting issue De Lange et al. (2021); Mai et al. (2022). It helps in efficiently leveraging existing knowledge and incorporating new information without the need for extensive retraining. The primary goal of CL is to develop techniques and strategies that allow a neural network to learn new tasks while retaining knowledge of previous tasks. In other words, it aims to enable the network to continually adapt to new information without completely erasing or degrading its performance on earlier tasks. Overall, CL helps to address the issue of catastrophic forgetting and minimizes the need for additional resources to store historical data. CL offers a range of strategies and methods, such as regularization (constraining weight update to avoid forgetting learned concepts), rehearsal (partially using some form of old data to replay with current data), and architectural modifications (reserving or partitioning network for different tasks), to help neural networks remember and consolidate knowledge from past tasks. These strategies help prevent or reduce catastrophic forgetting and improve the generalization ability of the model.